OpenStudy (anonymous):

Someone please help Elimination { x + y – 2z = 8 {5x – 3y + z =-6 {-2x – y + 4z = -13 Substitution {3x + y = 2z =22 {x + 5y + z =4 {x = -3z + 6

OpenStudy (igreen):

I can solve the second one for you..but the first equation is messed up. You put: 3x + y = 2z =22 There are 2 equal signs.

OpenStudy (igreen):

Can you fix that? @-EtherealEbullience

OpenStudy (anonymous):

{3x + y - 2z =22 {x + 5y + z =4 {x = -3z + 6

OpenStudy (igreen):

Okay, we need to change the first equation. 3x + y - 2z = 22 Add 2z to both sides: 3x + y = 2z + 22 Subtract 3x to both sides: y = -3x + 2z + 22

OpenStudy (texaschic101):

elimination : x + y - 2z = 8 -2x - y + 4z = -13 -----------------add -x + 2z = - 5 x + y - 2z = 8 ---->(3) 5x - 3y + z = -6 --------------- 3x + 3y - 6z = 24 (result of multiplying by 3) 5x - 3y + z = -6 --------------add 8x - 5z = 18 -x + 2z = -5 --->(8) 8x - 5z = 18 ----------- -8x + 16z = -40 (result of multiplying by 8) 8x - 5z = 18 -----------add 11z = - 22 z = -2 -x + 2z = -5 -x + 2(-2) = -5 -x - 4 = -5 -x = -5 + 4 -x = -1 x = 1 x + y - 2z = 8 1 + y - 2(-2) = 8 1 + y + 4 = 8 5 + y = 8 y = 8 - 5 y = 3 solution is : x = 1, y = 3, and z = -2

OpenStudy (anonymous):

Gosh. . .That's a long process. /.\